OpenStudy (anonymous):

How do you find the solutions of 5root3tanx+3=8root3tanx?

OpenStudy (anonymous):

\[5\sqrt{3}\tan x + 3 = 8\sqrt{3}\tan x\] \[3 = 3\sqrt{3}\tan x\]

OpenStudy (anonymous):

solve for tan x